On June 30, 2018, Campbell Company’s total current assets were $503,500 and its total current liabilities were $270,500. On July 1, 2018, Campbell issued a short-term note to a bank for $40,600 cash.

Required

  1. Compute Campbell’s working capital before and after issuing the note.

  2. Compute Campbell’s current ratio before and after issuing the note. (Round your answers to 2 decimal places.)

Before the transaction After the transaction
a. Working Capital
b. Current Ratio

On June 30, 2018, Solomon Company’s total current assets were $504,000 and its total current liabilities were $275,000. On July 1, 2018, Solomon issued a long-term note to a bank for $38,400 cash.

Required

  1. Compute Solomon’s working capital before and after issuing the note.

  2. Compute Solomon’s current ratio before and after issuing the note. (Round your answers to 1 decimal place.)

Before the transaction After the transaction
a. Working Capital
b. Current Ratio
Answer #1

Calculate following

Before the transaction After the transaction
a Working capital 503500-270500 = 233000 (503500-40600)-(270500-40600) = 233000
b Current ratio 503500/270500 = 1.9 462900/229900 = 2.0